Question 2 (20 marks)

Sunshine Corporation purchased a new machine on Jan 4, 2018 for $185,000 cash. The machine has a useful life of five years or of 26,000 hours. After the useful life the machine will have a residual value of $3,000. The machine was used for 3,600 hours in 2018 and 4,500 hours in 2019.

Required:

  1. Calculate the depreciation expense for 2018 and 2019 under each of the following methods:
  1. Straight-line (3 marks)
  2. Double diminishing-balance (5 marks)
  3. Units-of -production (5 marks)

  1. Calculate the carrying amount at the end of 2019 for the machine when the company is using the Straight-line depreciation method. (3 marks)

  1. Record the journal entry for the cash purchase of the machine on Jan 4, 2018 and for its depreciation expense for the year ended December 31, 2019 under the units-of-production method. (4 marks)
Answer #1

Part a)Calculation of Depreciation expense for 2018 and 2019

i) Straight Line Method

Depreciation =(Cost - Salvage value)/Useful Life

Depreciation each year =(185000 -3000)/5=$36400

Depreciation for 2018 =$36400

Depreciation for 2019 =$36400

ii)Double Diminishing Balance

Depreciation Rate under SLM=36400/182000*100 =20%

Depreciation rate under DDBM =20*2 =40%

Year Beginning value Depreciation expense Ending Value
2018 185000 185000*40%=74000 111000
2019 111000 111000*40%=44400 66600


Iii) Units of Production Method

Depreciation Rate per hour =(Cost -Salvage value)/useful hours

=(185000-3000)/26000 =$7 per hour

Depreciation for year 2018 =3600 hours *7 =$25200

Depreciation for year 2019 = 4500 hours *7 =$31500

part b) Carrying amount at end of 2019 under SLM

carrying amount (2019)= Book value - Depreciation for 2018 and 2019

Carrying Amount(2019)=185000 -36400 -36400 =$112200

part c) Journal entry


Date Particulars Debit Credit
Jan 4,2018

Machine

To cash

( Being machine purchased in cash)

185000 185000
Dec 31, 2019

Depreciation

To Accumulated Depreciation

( Being depreciation charged for 2019 under units of production method)